# Dark Matter Mounds: towards a realistic description of dark matter overdensities around black holes

### v1

@@ Abstract
Figure 1: Density profile of a $10^{5}\,M_{\odot}$ polytropic Supermassive Star (SMS) with polytropic index $n=3$, compared to the density profile of the dark matter halo hosting it, with virial mass $M_{\rm vir}=10^{7}\,M_{\odot}$ and concentration $c=3$ at redshift $z_{\rm f}=15$. The yellow line corresponds to the spike solution of Ref. [4].
Supermassive star formation. We assume a set of fiducial values for the parameters describing the supermassive star (SMS) and DM density profile, which are motivated by the results of recent cosmological simulations of radiation-driven direct-collapse BH formation [39]. Our formalism can be easily generalised to other scenarios, as long as the SMS forms at the center of the DM halo. We consider in particular an SMS with mass $M_{\rm SMS}=10^{5}M_{\odot}$, with central density $\rho_{c}=1$ g cm-3 [34]. The interior of an SMS is supported by radiation pressure, and while detailed calculations of stellar evolution show that an accreting SMS has a core-envelope structure with a bloated, high-entropy atmosphere [34, 40], the core structure is well approximated by an $n=3$ polytrope.
@@
Adiabatic growth. In order to compute the dark matter density after SMS formation, we need to generalise the GS spike formation formalism. Assuming spherical symmetry, the angular momentum $L$ and the radial action $I(\mathcal{E},L)$ are adiabatic invariants. We denote with $\cal{E}$ the relative energy per unit mass ${\cal E}=\Psi(r)-v^{2}/2$, where $\Psi(r)=\phi_{0}-\phi(r)$ is the positive-definite relative gravitational potential, so that all particles with ${\cal E}>0$ are on bound orbits. We can write the radial action as
@@
Collapse. The accreting supermassive star collapses to a BH after about $10^{5}$ year. The exact timing of the gravitational collapse depends on the gas mass accretion rate [40, 45], but it occurs roughly near the end of the hydrogen burning phase. After the onset of implosion, either after hydrogen exhaustion or when triggered by general relativistic instability, the star contracts and collapses very rapidly. The sudden change in gravitational potential modifies the orbits of individual dark matter particles. The angular momentum of the initial and final orbits will be the same, but the final energy $\mathcal{E}_{f}$ is related to the initial energy $\mathcal{E}_{i}$ as $\mathcal{E}_{f}=\mathcal{E}_{i}+\Delta\Psi(r)$, where $\Delta\Psi$ is the difference between the final potential due to the BH ($m_{\mathrm{BH}}=10^{5}\,M_{\odot}$) and the initial potential due to the supermassive star.
Figure 2: Top panel: Density profile of dark matter after the formation of the supermassive star (red), and after the collapse to a black hole (blue). For comparison, we show the spike solution of Ref. [4] (yellow). The vertical lines denote the cutoff radii $2r_{s}$, the separation of the EMRI binary 4 years before merger, and the radius of the SMS. Bottom panel: Logarithmic slope of the profiles. The horizontal lines correspond to the slopes $-7/3$ and $-1$, see text for further details.
@@
$P(r|\mathcal{E},L)\,\mathrm{d}r=\frac{2}{T(\mathcal{E},L)}\frac{1}{v_{r}(r,\mathcal{E},L)}\,\mathrm{d}r\,,$ (4)
where $T(\mathcal{E},L)=2\int_{r_{\mathrm{min}}}^{r_{\mathrm{max}}}v_{r}^{-1}\,\mathrm{d}r$ is the period of the orbit. Ref. [23] computed the final density profile in the case of initially circular orbits, whereas we are interested here in the more complex case of an initial DM distribution arising from the adiabatic contraction of an initial NFW profile following the SMS growth. To obtain the final density profile, we thus need to integrate the probability in Eq. 4 over the appropriate initial distribution function:
$\rho_{f}(r)=\frac{1}{4\pi r^{2}}\int\mathrm{d}^{3}\mathbf{r}_{i}\,\mathrm{d}^{3}\mathbf{v}_{i}\frac{2f_{i}(\mathcal{E}_{i},L)}{T(\mathcal{E}_{f},L)\,v_{r}(r,\mathcal{E}_{f},L)}\,,$ (5)
@@
$\rho_{f}(r)=\frac{8\pi}{r^{2}}\int\,\mathrm{d}r_{i}\,\frac{L\,\mathrm{d}L\,\mathrm{d}{\cal E}_{i}}{v_{r}(r_{i},\mathcal{E}_{i},L)\,}\frac{f_{i}(\mathcal{E}_{i},L)}{T(\mathcal{E}_{f},L)\,v_{r}(r,\mathcal{E}_{f},L)}\,.$ (6)
@@
$\rho_{f}(r)=\frac{8\pi}{r^{2}}\times\frac{1}{N}\sum_{k}\frac{w_{k}}{T(\mathcal{E}_{f,k},L_{k})\,v_{r}(r,\mathcal{E}_{f,k},L_{k})}\,.$ (7)
In order to account for capture by the black hole, we exclude from the sum all particles on final orbits with $r_{\mathrm{min}}\leq 2r_{s}=4G_{N}m_{\mathrm{BH}}/c^{2}$, accounting for relativistic effects following Ref. [46].
@@
Results. In Fig. 2, we show the dark matter density profile after the formation of the supermassive star (solid red), and after the collapse to a black hole (solid blue). For comparison, we also show the spike profile expected from the adiabatic growth of a $10^{5}\,M_{\odot}$ BH (“GS profile”), without the intermediate stage of SMS formation. The grey shaded region in Fig. 2 highlights the range of radii probed by an extreme mass-ratio inspiral (EMRI, $m_{1}=10^{5}\,M_{\odot}$, $m_{2}=10\,M_{\odot}$), starting 4 years before the merger.
@@
We see that the density after SMS formation is significantly lower than the spike profile of Ref. [4], in the region where the gravitational wave signal is expected to peak. Note also that the density profile extends below the Schwarzschild radius of the black hole, since collapse has not taken place yet. When the SMS collapses to a BH, the profile steepens again, due to the increased potential experienced by particles inside the star, but remains below the spike solution, except in a small range of radii ($r\sim 10^{-6}\,\mathrm{pc}$), where the profile slightly exceeds the standard spike solution. This ‘excess’ can be understood in terms of mass conservation. The central density is lower than in the standard spike solution (because adiabatic contraction is less efficient at $r\ll r_{\mathrm{SMS}}$), meaning that the density at radii just below $r\sim r_{\mathrm{SMS}}$ must be larger.
@@
In the lower panel of Fig. 2, we show the logarithmic slope of the DM density profiles at each stage. Overall, following the evolution of the dark matter density profile through the formation of the SMS, and then through its collapse, leads to a profile which matches the spike solution ($\gamma=-7/3$) outside the radius of the SMS but is significantly shallower inside. The behaviour of the mounds can be understood qualitatively, as we demonstrate for circular orbits in the Supplementary Material. Conservation of DM mass and angular momentum in the initial phase of adiabatic growth of the SMS leads to a slope shallower than -1, while the subsequent collapse to a BH should lead to a slope of -5/4, further flattened to a slope of around $\gamma=-1$ by the effects of capture by the BH.
@@
In Fig. 3, we show the DM density profile following the subsequent adiabatic growth of the DCBH from $10^{5}\,M_{\odot}$ to $2\times 10^{5}\,M_{\odot}$ (light purple) and to $10^{6}\,M_{\odot}$ (dark purple). We also show a comparison with the GS profile for these two final BH masses, which includes a cut-off at $4r_{s}$, proportional to $(1-4r_{s}/r)^{3}$ [4] . As we increase the final BH mass, we find that the spike normalisation at large radii grows as $m_{\mathrm{BH}}^{2/3}$, as expected from the GS formalism, up to a factor of $\sim 5$ (for growth up to $10^{6}\,M_{\odot}$). Increasing the final mass also increases the Schwarzschild radius of the BH, leading to a cut-off in the spike at larger and larger radii. With growth by a factor of 2, the flattened inner density profile is still visible in the spike. For growth by a factor of 10, however, the outer power-law slope appears to persist down to the cut-off. In this case, the final DM density exceeds the spike prediction by more than an order of magnitude at small radii. While the density of the post-collapse mound in Fig. 2 is smaller than the spike prediction at $r\sim 2\times 10^{-7}\,\mathrm{pc}$, it exceeds the spike prediction at intermediate radii ($r\sim 10^{-6}\,\mathrm{pc}$), as a direct consequence of the flattening of the density profile. This excess is then amplified by subsequent adiabatic growth.
@@
Figure 3: Density profile of dark matter after the formation of the direct collapse black hole ($m_{\mathrm{BH}}=10^{5}\,M_{\odot}$, blue line) and after the subsequent adiabatic growth of the BH by a factor of 2 (light purple) and a factor of 10 (dark purple). For comparison, we show the spike profile from Ref. [4] (dashed lines), assuming the adiabatic growth of a BH with the given mass from a seed of negligible mass at the centre of an NFW halo, without the intermediate stages described above.
@@
In order to compute the evolution of dark matter after collapse, we have introduced a new method, based on a Monte Carlo procedure, which generalises the results of Ref. [23] to non-circular orbits, and includes the relativistic corrections to the inner cut-off from Ref. [46]. We have demonstrated that the DM profile around the BH is significantly shallower than previously thought on scales smaller than the SMS radius, right after collapse (Fig. 2, bottom panel), leading to the formation of rounded peaks that we have dubbed ‘mounds’ to discriminate them from the sharper ‘spikes’ introduced in Ref. [4]. If the BH grows substantially in mass from this initial seed, the combined effect of instantaneous collapse of the SMS to a BH, and the inclusion of relativistic effects, leads to a final profile that is similar to a ‘spike’, but which extends further in, and reaches DM densities up to two orders of magnitude larger than previously thought.
@@
Appendix A Supplementary Material: Analytic Arguments
@@
$\begin{split}\rho_{i}(r_{i})\,r_{i}^{2}\,\mathrm{d}r_{i}&\sim\rho_{f}(r_{f})\,r_{f}^{2}\,\mathrm{d}r_{f}\\ \Rightarrow r_{i}^{3+\gamma_{i}}&\sim r_{f}^{3+\gamma_{f}}\,,\end{split}$ (8)
@@
Figure 4: Same as Fig. 2, but comparing the density profile after DCBH formation with (solid blue) and without (dashed blue) accounting for the capture of orbits by the newly formed BH. Without accounting for capture, the inner logarithmic slope of the DM density profile is $-5/4$, in agreement with our analytic arguments. This is then flattened towards $-1$ when including the effects of capture.
@@
$\begin{split}v_{r}(r,r_{0})=&\sqrt{\frac{2GM_{\mathrm{BH}}}{rr_{0}}}\\ &\times\sqrt{\left(r_{0}-r\right)\left(1-\frac{m_{\mathrm{SMS}}\left(r_{0}\right)}{2m_{\mathrm{BH}}}\left(1+\frac{r_{0}}{r}\right)\right)}\,,\end{split}$ (12)
@@
where we have defined the dimensionless quantities $x=r/r_{\mathrm{max}}$ and $x_{0}=r_{0}/r_{\mathrm{max}}$. For small $r$, the lower limit tends to 00 and the dimensionless integral becomes roughly independent of $r$. Recalling that $r_{\mathrm{max}}\sim r^{1/4}$ and that $\gamma_{i}\approx-0.5$, we thus expect $\rho_{f}(r)\sim r^{-5/4}$ after collapse.
In Fig. 4, we show the DM density profile after DCBH formation which we obtain in our full numerical analysis, when the effects of capture by the BH are ignored (dashed blue line). As we see in the lower panel, the inner density profile in this case matches very well the slope of $-5/4$ derived above. This power-law is further flattened once we include the effects of capture (solid blue line). We assume that a DM particle will be captured if the pericentre $r_{\mathrm{peri}}$ of its orbit is less than $r_{\mathrm{cap}}=2r_{s}$. This requirement equates roughly to requiring $r_{0}>(2r_{\mathrm{cap}}r_{\mathrm{SMS}}^{3})^{1/4}$, thus altering the lower limit of the integral in Eq. 13, from $x$ to $(2r_{\mathrm{cap}}r_{\mathrm{SMS}}^{3})^{1/4}/r_{\mathrm{max}}\sim(r_{\mathrm{cap}}/r)^{1/4}$. The lower limit of the dimensionless integral now grows as $r\rightarrow 0$ and the integral now scales roughly as $(1-(r_{\mathrm{cap}}/r)^{1/4})$ for sufficiently small $r$. This leads to a smooth flattening from a slope of $-5/4$ towards zero (eventually turning over and giving a cut-off at $r=r_{\mathrm{cap}}$). The behaviour shown in Fig. 2 is more complicated; in that case, the initial velocities are not circular and so the behaviour of $r_{\mathrm{peri}}$ with initial radius is more complicated.

### v2

# myurlcolorviolet

@@ Abstract
Figure 1: Density profile of a $10^{5}\,M_{\odot}$ polytropic Supermassive Star (SMS) with polytropic index $n=3$, compared to the density profile of the dark matter halo hosting it, with virial mass $M_{\rm vir}=10^{7}\,M_{\odot}$ and concentration $c=3$ at redshift $z_{\rm f}=15$. The yellow lines corresponds to the non-relativistic spike profile of Ref. [4] and the relativistic profile of Ref. [39].
Supermassive star formation. We assume a set of fiducial values for the parameters describing the supermassive star (SMS) and DM density profile, which are motivated by the results of recent cosmological simulations of radiation-driven direct-collapse BH formation [40]. Our formalism can be easily generalised to other scenarios, as long as the SMS forms at the center of the DM halo. We consider in particular a spherically symmetric SMS with mass $M_{\rm SMS}=10^{5}M_{\odot}$, with central density $\rho_{c}=1$ g cm-3 [34]. The interior of an SMS is supported by radiation pressure, and while detailed calculations of stellar evolution show that an accreting SMS has a core-envelope structure with a bloated, high-entropy atmosphere [34, 41], the core structure is well approximated by an $n=3$ polytrope.
@@
[…]
@@
Collapse. The accreting supermassive star collapses to a BH after about $10^{5}$ year. The exact timing of the gravitational collapse depends on the gas mass accretion rate [41, 46], but it occurs roughly near the end of the hydrogen burning phase. After the onset of implosion, either after hydrogen exhaustion or when triggered by general relativistic instability, the star contracts and collapses very rapidly. We assume that the entire mass of the SMS collapses to form the BH. The sudden change in gravitational potential modifies the orbits of individual dark matter particles. The angular momentum of the initial and final orbits will be the same, but the final energy $\mathcal{E}_{f}$ is related to the initial energy $\mathcal{E}_{i}$ as $\mathcal{E}_{f}=\mathcal{E}_{i}+\Delta\Psi(r)$, where $\Delta\Psi$ is the difference between the final potential due to the BH ($m_{\mathrm{BH}}=10^{5}\,M_{\odot}$) and the initial potential due to the supermassive star.
Figure 2: Top panel: Density profile of dark matter after the formation of the supermassive star (red), and after the collapse to a black hole (blue). For comparison, we show the relativistic spike solution of Ref. [39] (yellow). The vertical lines denote the cutoff radii $2r_{s}$, the separation of the EMRI binary 4 years before merger, and the radius of the SMS. Bottom panel: Logarithmic slope of the profiles. The horizontal lines correspond to the slopes 7/3-7/3 and 1-1, see text for further details.
@@
$P(r|\mathcal{E},L)\,\mathrm{d}r=\frac{2}{T(\mathcal{E},L)}\frac{1}{\tilde{v}_{r}(r,\mathcal{E},L)}\,\mathrm{d}r\,$, (4)
where $T(\mathcal{E},L)=2\int_{r_{\mathrm{min}}}^{r_{\mathrm{max}}}\tilde{v}_{r}^{-1}\,\mathrm{d}r$ is the period of the orbit. Post-collapse, we use the radial velocity in the Schwarzschild metric [39]:
$\tilde{v}_{r}^{2}=c^{2}-2\mathcal{E}-\left(c^{2}-\frac{2GM_{\mathrm{BH}}}{r}\right)\left(1+\frac{L^{2}}{r^{2}c^{2}}\right)\,.$ (5)
Ref. [23] computed the final density profile in the case of initially circular orbits, whereas we are interested here in the more complex case of an initial DM distribution arising from the adiabatic contraction of an initial NFW profile following the SMS growth. To obtain the final density profile, we thus need to integrate the probability in Eq. 4 over the appropriate initial distribution function:
$\rho_{f}(r)=\frac{1}{4\pi r^{2}}\int\mathrm{d}^{3}\mathbf{r}_{i}\,\mathrm{d}^{3}\mathbf{v}_{i}\frac{2f_{i}(\mathcal{E}_{i},L)}{T(\mathcal{E}_{f},L)\,\tilde{v}_{r}(r,\mathcal{E}_{f},L)}\,$, (6)
@@
$\rho_{f}(r)=\frac{8\pi}{r^{2}}\int\,\mathrm{d}r_{i}\,\frac{L\,\mathrm{d}L\,\mathrm{d}{\cal E}_{i}}{v_{r}(r_{i},\mathcal{E}_{i},L)\,}\frac{f_{i}(\mathcal{E}_{i},L)}{T(\mathcal{E}_{f},L)\,\tilde{v}_{r}(r,\mathcal{E}_{f},L)}\,.$ (7)
@@
$\rho_{f}(r)=\frac{8\pi}{r^{2}}\times\frac{1}{N}\sum_{k}\frac{w_{k}}{T(\mathcal{E}_{f,k},L_{k})\,\tilde{v}_{r}(r,\mathcal{E}_{f,k},L_{k})}\,.$ (8)
We apply cuts on $(\mathcal{E}_{f},L)$ following the prescription in Ref. [39], in order to remove orbits which are unstable or are immediately captured by the BH. Full details of these cuts (and a comparison with the fully relativistic calculation) are given in the Supplementary Material.
@@
Results. In Fig. 2, we show the dark matter density profile after the formation of the supermassive star (solid red), and after the collapse to a black hole (solid blue). For comparison, we also show the spike profile expected from the adiabatic growth of a $10^{5}\,M_{\odot}$ BH (“Sadeghian et al. Spike”), without the intermediate stage of SMS formation. The grey shaded region in Fig. 2 highlights the range of radii probed by an extreme mass-ratio inspiral (EMRI, $m_{1}=10^{5}\,M_{\odot}$, $m_{2}=10\,M_{\odot}$), starting 4 years before the merger.
@@
We see that the density after SMS formation is significantly lower than the spike profile of Refs. [4, 39], in the region where the gravitational wave signal is expected to peak. Note also that the density profile extends below the Schwarzschild radius of the black hole, since collapse has not taken place yet. When the SMS collapses to a BH, the profile steepens again, due to the increased potential experienced by particles inside the star, but remains below the spike solution, except in a small range of radii ($r\sim 10^{-6}\,\mathrm{pc}$), where the profile slightly exceeds the standard spike solution. This ‘excess’ can be understood in terms of mass conservation. The central density is lower than in the standard spike solution (because adiabatic contraction is less efficient at $r\ll r_{\mathrm{SMS}}$), meaning that the density at radii just below $r\sim r_{\mathrm{SMS}}$ must be larger.
@@
In the lower panel of Fig. 2, we show the logarithmic slope of the DM density profiles at each stage. Overall, following the evolution of the dark matter density profile through the formation of the SMS, and then through its collapse, leads to a profile which matches the spike solution ($\gamma=-7/3$) outside the radius of the SMS but is significantly shallower inside. We find this behavior for a wide range of values for the total DM mass and the SMS mass. For more massive SMSs, we find that the shallower inner region is larger, as expected. The behaviour of the mounds can be understood qualitatively, as we demonstrate for circular orbits in the Supplementary Material. Conservation of DM mass and angular momentum in the initial phase of adiabatic growth of the SMS leads to a slope shallower than 1-1, while the subsequent collapse to a BH should lead to a slope of 5/4-5/4, further flattened to a slope of around $\gamma=-1$ by the effects of capture by the BH.
@@
In Fig. 3, we show the DM density profile following the subsequent adiabatic growth of the DCBH from $10^{5}\,M_{\odot}$ to $2\times 10^{5}\,M_{\odot}$ (light purple) and to $10^{6}\,M_{\odot}$ (dark purple). We also show a comparison with the Sadeghian et al. relativistic spike profile for these two final BH masses [39]. As we increase the final BH mass, we find that the spike normalisation at large radii grows as $m_{\mathrm{BH}}^{2/3}$, as expected from the GS formalism, up to a factor of $\sim 5$ (for growth up to $10^{6}\,M_{\odot}$). Increasing the final mass also increases the Schwarzschild radius of the BH, leading to a cut-off in the spike at larger and larger radii. With growth by a factor of 2, the flattened inner density profile is still visible in the spike. For growth by a factor of 10, however, the outer power-law slope appears to persist down to the cut-off. In this case, the final DM density matches that expected for the adiabatic growth of an infinitesimal seed down to a few Schwarzschild radii, with a mild suppression of the density below this radius.
@@
Figure 3: Density profile of dark matter after the formation of the direct collapse black hole ($m_{\mathrm{BH}}=10^{5}\,M_{\odot}$, blue line) and after the subsequent adiabatic growth of the BH by a factor of 2 (light purple) and a factor of 10 (dark purple). For comparison, we show the corresponding relativistic spike profile following the formalism of Ref. [39] (dashed lines), assuming the adiabatic growth of a BH with the given mass from a seed of negligible mass at the centre of an NFW halo, without the intermediate stages described above.
@@
In order to compute the evolution of dark matter after collapse, we have introduced a new method, based on a Monte Carlo procedure, which generalises the results of Ref. [23] to non-circular orbits, and includes the relativistic corrections to the inner cut-off from Ref. [39]. We have demonstrated that the DM profile around the BH is significantly shallower than previously thought on scales smaller than the SMS radius, right after collapse (Fig. 2, bottom panel), leading to the formation of rounded peaks that we have dubbed ‘mounds’ to discriminate them from the sharper ‘spikes’ introduced in Ref. [4]. If the BH grows substantially in mass from this initial seed, the combined effect of instantaneous collapse of the SMS to a BH, and the inclusion of relativistic effects, leads to a final profile that is similar to a ‘spike’, but which may be flatter in the innermost regions.
@@
Appendix A Supplementary Material
A.1 Relativistic Density Profile
Here, we present further details of how we incorporate relativistic effects in the analysis. While the energy before and after adiabatic growth are calculated non-relativistically, we include relevant relativistic effects in the final density profile through (a) the use of the radial velocity in the Schwarzschild metric in Eq. (5); and (b) cuts on the energy $\mathcal{E}$ and angular momentum $L$ to account for DM capture by the black hole.
These cuts are described in Ref. [39] and exclude orbits which pass through the Schwarzschild radius, as well as unstable orbits which will eventually be captured. To first order in $\mathcal{E}$, the relativistic energy of the particles per unit mass can be written as $\mathcal{E}_{\mathrm{rel}}\approx c^{2}$-$\mathcal{E}$. With this, the cuts amount to removing particles at a radius $r$ with:
$\begin{split}L<L_{c}&\equiv 2r_{s}c\left(1-\frac{4\mathcal{E}}{c^{2}}\right)^{1/2}\\ \mathcal{E}>\mathcal{E}_{c}&\equiv\frac{c^{2}}{2}\frac{r_{s}(r-2r_{s})}{(r^{2}+2rr_{s}-4r_{s}^{2})}\,.\end{split}$ (9)
This correctly gives rise to a cut-off in the spike density profile at $r=2r_{s}$ as discussed in Ref. [39].
In order to verify that this prescription provides an accurate estimate of the density profile, we show in Fig. 4 the non-relativistic spike of Gondolo & Silk [4] (dotted line); the fully relativistic spike of Sadeghian et al. [39] (dashed line); and the formalism including relativistic corrections we present here (solid line). In all cases, the spike is grown adiabatically from an infinitesimal seed (without the intermediate stages described in the main text). We see that our formalism matches very closely that of Sadeghian et al. down to $r\lesssim 10r_{s}$. Below this radius, our result deviates by at most 30% from the fully-relativistic result, though it provide a substantial improvement over the non-relativistic GS spike. We are therefore confident that the most important relativistic effects are included in the analysis, especially considering at below $\sim 5\,r_{s}$ the dynamics of IMRIs/EMRIs is expected to be dominated by GW emission, for which the precise DM spike density will be less relevant.
Figure 4: Comparison of our formalism for adiabatic growth (from an infinitesimal seed) with the full relativistic formalism of Sadeghian et al. [39] and the non-relativistic formalism of Gondolo & Silk [4]. Here, we assume the adiabatic growth of a BH of mass $1.1\times 10^{6}\,M_{\odot}$ at the centre of a Hernquist DM profile, as used in Ref. [39]. The dotted and dashed lines are taken from Fig. 3 of Ref. [39].
A.2 Analytic Arguments
@@
[…]
@@
Figure 5: Same as Fig. 2, but comparing the density profile after DCBH formation with (solid blue) and without (dashed blue) accounting for the capture of orbits by the newly formed BH. Without accounting for capture, the inner logarithmic slope of the DM density profile is 5/4-5/4, in agreement with our analytic arguments. This is then flattened towards 1-1 when including the effects of capture.
@@
[…]
@@
Figure 6: The top 3 panels show the density profile after SMS formation, and after DCBH formation, for a halo mass of $10^{8}$ $M_{\odot}$, and $M_{\mathrm{SMS}}$ equal to $10^{5}$ (left), $10^{6}$ (center) and to $10^{7}$ $M_{\odot}$ (right). The bottom panel shows the corresponding logarithmic slope as a function of radius.
@@
where we have defined the dimensionless quantities $x=r/r_{\mathrm{max}}$ and $x_{0}=r_{0}/r_{\mathrm{max}}$. For small $r$, the lower limit tends to 0 and the dimensionless integral becomes roughly independent of $r$. Recalling that $r_{\mathrm{max}}\sim r^{1/4}$ and that $\gamma_{i}\approx-0.5$, we thus expect $\rho_{f}(r)\sim r^{-5/4}$ after collapse.
In Fig. 5, we show the DM density profile after DCBH formation which we obtain in our full numerical analysis, when the effects of capture by the BH are ignored (dashed blue line). As we see in the lower panel, the inner density profile in this case matches very well the slope of 5/4-5/4 derived above. This power-law is further flattened once we include the effects of capture (solid blue line). We assume that a DM particle will be captured if the pericentre $r_{\mathrm{peri}}$ of its orbit is less than $r_{\mathrm{cap}}=2r_{s}$. This requirement equates roughly to requiring $r_{0}>(2r_{\mathrm{cap}}r_{\mathrm{SMS}}^{3})^{1/4}$, thus altering the lower limit of the integral in Eq. 15, from $x$ to $(2r_{\mathrm{cap}}r_{\mathrm{SMS}}^{3})^{1/4}/r_{\mathrm{max}}\sim(r_{\mathrm{cap}}/r)^{1/4}$. The lower limit of the dimensionless integral now grows as $r\rightarrow 0$ and the integral now scales roughly as $(1-(r_{\mathrm{cap}}/r)^{1/4})$ for sufficiently small $r$. This leads to a smooth flattening from a slope of 5/4-5/4 towards zero (eventually turning over and giving a cut-off at $r=r_{\mathrm{cap}}$). The behaviour shown in Fig. 2 is more complicated; in that case, the initial velocities are not circular and so the behaviour of $r_{\mathrm{peri}}$ with initial radius is more complicated.
A.3 Robustness tests
In this section, we demonstrate the robustness of our results against variations in halo mass and SMS mass values. We show the results for two different scenarios:
  • We keep the SMS mass fixed to a value of $10^{5}$ $M_{\odot}$ and vary the halo mass from $10^{6}$ to $10^{8}$ $M_{\odot}$.
  • We keep the halo fixed to a value of $10^{8}$ $M_{\odot}$ and vary $M_{\mathrm{SMS}}$ from $10^{5}$ to $10^{7}$ $M_{\odot}$.
In all cases we keep the concentration constant ($c=3$), given the negligible variation of this parameter with the halo mass at the redshift of interest [43].
We find very little dependence of our results on the halo mass, as expected given that the region that is relevant for our treatment is dominated by the potential of the central object. Regarding variations of the SMS mass, we observe an enlargement of the flatter inner region for larger values of this parameter, as expected. We show in Fig. 6 a set of plots that illustrate this result, showing both the change in the profile and the change in the behavior of the log slope as a function of radius, for 3 different values of the SMS mass.